Please help me solve this equation:
log(x+3)=2+log(x)

Answer by stanbon(75887)

Please help me solve this equation:
log(x+3)=2+log(x)
----
log(x+3) - log(x) = 2
----
log[(x+3)/x] = 2
----
(x+3)/x = 10^2
---
x+3 = 100x
99x = 3
x = 1/33
===============
Cheers,
Stan H.
=====================